# COOL-LAMPS. VII. Quantifying Strong-lens Scaling Relations with 177 Cluster-scale Gravitational Lenses in DECaLS

### v3

@@ 2.4 Photometry
We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$, $r$, and $z$-band imaging data from DECaLS LS DR9. Since it is critical that we do not include light from the lensed source arcs as well as any intervening stars or foreground galaxies in the measured photometry, we created two masks to exclude non-lens-galaxy regions. The first mask removed flux from the lensed source arcs, and the second mask removed flux from interlopers as identified by a color and magnitude screening by eye. Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1, and the resulting mean and standard deviation were adopted as the $g$, $r$, and $z$-band photometry and their errors respectively for each system.
@@ 3.1 Total Mass
Following Remolina González et al. (2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2)—where $D(z)$ refers to the angular diameter distance at redshift $z$—via a Monte Carlo approach. The enclosed total mass was computed 1000 times for each system, in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with mean equal to the parameter value and standard deviation equal to the parameter error. Also described in Remolina González et al. (2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed source arcs in each system—where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al. (2020). This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system, in which larger-Einstein-radius systems have a stronger offset. The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly symmetric systems.
@@
For the purposes of applying this correction in each of the 177 lensing systems analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system. For example, a lens with a tangential arc stretching from an azimuthal angle of 10 to 100 degrees would have $\phi=0.25$. As was done by Remolina González et al. (2020), $\phi$ was used as an observable proxy for the degree of symmetry in each system to determine whether or not the correction factor was needed. Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark-matter halo and thus require the empirical correction. For spherically symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed. We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the the regions masking the lensed source arcs described in Section 2.4 in each system. Adopting the convention used by Remolina González et al. (2020) in their analysis, if $\phi<0.5$, Equation (3) was applied. If $\phi\geq 0.5$, it was not applied. Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error respectively for each system.
@@ 3.2 Stellar Luminosity and Mass
We conducted parametric SED fitting using the photometry obtained in Section 2.4 with Prospector (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2021, 2023); five free parameters and three assumed parameters parameterized each SED. A summary of all eight parameters used in fitting can be found in Table 1. Utilizing emcee (Foreman-Mackey et al., 2013) as implemented in Prospector with 84 walkers for a total of 6720 iterations, only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) were taken as representing the posterior distribution for each of the five free parameters in each system in order to eliminate the significant burn-in sequence of the fitting process. After fitting, we adopted the 50th percentile of the posterior distribution as the value for each free parameter, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error. Observed-frame SEDs for the systems shown in Figure 1 after SED fitting are shown in Figure 2.
To derive luminosity, we generated new SEDs with a random sample of 1024 free-parameter vectors from the posterior distributions of the best-fit SED in each system. We integrated over the rest-frame wavelength interval of 3000Å to 7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts). This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths which are significantly more luminous in cluster galaxies. After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random parameter vectors as the luminosity and its error respectively for each system.
@@ 4 Results
As Figure 3 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated. Indeed, we note that the most massive clusters in this work also possess a large stellar mass and thus have a large luminosity. Using emcee, we constrained a plane-of-best-fit using a likelihood maximization estimation taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$—yielding:
$Z=0.137_{-0.051}^{+0.051}X+0.522_{-0.023}^{+0.022}Y+4.207_{-0.290}^{+0.300}$ (4)
This simple and analyic relationship derived from a sample of known strong lenses, considering only quantities derived from observations, may have utility in differentiating bona fide strong lenses from non-lenses in large samples of candidate lenses where the evidence of lensing is not nearly as robust as the sample of clusters analyzed here. Moreover, this correlation offers a future opportunity to test simulations against observations—particularly those that aim to generate realistic galaxy populations in dense cluster environments.
Figure 3: A 3D visualization of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles. A plane-of-best-fit is also shown in black as well. Errors in data on each axis are represented as red, blue, and green lines respectively.
@@ 5.1 Scaling Relations
Besides the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system using the stellar-mass posterior distribution from each SED constrained with Prospector. We obtained this by correcting the derived enclosed stellar mass formed in each system with the surviving mass fraction returned by Prospector. This is necessary because the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the galaxy (e.g., Li et al. 2017)—the latter being what Prospector parametrically fits. While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong-lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
@@ 5.1.1 Total Mass — Stellar Mass
Figure 4: The total mass enclosed within the Einstein aperture for each system on the x-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the y-axes. Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000Å to 7000Å. Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture. Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
@@
In Panel I of Figure 4, we see a clearly positive correlation between the enclosed total mass and the enclosed stellar mass for all systems. By performing a linear regression with emcee using a likelihood maximization estimation that accounts for errors in data on both axes, we find that the slope of this correlation is $0.443\pm 0.035$ from the posterior distribution of the MCMC fit. We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall intrinsic scatter. In Table 2 of Kravtsov et al. (2018), they found $M_{500}$ (White, 2001) and the stellar mass of solely the BCG to be correlated with a logarithmic slope of $0.33\pm 0.11$. In addition, Kravtsov et al. (2018) found $M_{500}$ and the stellar mass of the entire cluster to be correlated with a logarithmic slope of $0.59\pm 0.08$. The fact that the slope for this work falls between the two slopes quoted by Kravtsov et al. (2018) is consistent with how our methodology samples a physical regime between exclusively the BCG and that of the entire cluster.
Panel II of Figure 4 further explores this correlation, with only those clusters where multiple cluster members were contained within the Einstein aperture are plotted. A linear regression for the multiple-galaxy systems gives a slope of $0.453\pm 0.049$, which is consistent with the slope using all systems.
5.1.2 Total Mass — Stellar-to-Total Mass Fraction
@@
In Panel III of Figure 4, we plot the enclosed stellar-to-total mass fraction against the enclosed total mass and see a clearly negative correlation (again with the same colormapping as Panel I to emphasize intrinsic scatter). Also using emcee, we find that the slope of this correlation is $-0.563\pm 0.035$. Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.561\pm 0.050$. We again find that the slope for the multiple-galaxy systems are statistically indistinguishable from the slope for the entire sample shown in Panel III. Andreon (2010) measured 52 clusters over a slightly larger mass range and found that the stellar mass fraction depended on halo mass with a slope of $-0.55\pm 0.08$ when measuring out to a physical extent of $R_{200}$. Both of the slopes in Panels III and IV are statistically similar to Andreon (2010), and while the linear fit derived by Andreon (2010) has a slightly larger normalization, this is to be expected because the cluster masses are constrained out to $R_{200}$.
@@ 5.2 Photometric Bias and Error
Deriving light in Section 2.4 via directly integrating over the Einstein aperture presents some challenges. Since we identified red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically bluer galaxies at redshift $z>z_{{}_{BCG}}$ that appear to be the same color as the cluster members. Conversely, we may have excluded intrinsically bluer bona fide cluster members. However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al. 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light. In addition, bluer galaxies at redshift $z>z_{{}_{BCG}}$ appear relatively fainter, which are also unlikely to bias our methodology of measuring cluster light. Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying lens-galaxy flux is typically minimal due to the limited angular size of such stars.
@@
We also note that our sample contains extremely few examples of systems which have a measured Einstein radius of $\theta_{E}\lessapprox 2^{\prime\prime}$. To some extent, this is a consequence of our chosen sample of cluster-scale lenses, which we would generally expect to have larger Einstein radii. However, we caution that there exists a large scatter between the total halo mass and Einstein radius (Fox et al., 2022). Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernable BCG light. This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light. While we thus expect this work to undersample the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$, $r$, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0 respectively in the average case of two exposures per filter (Dey et al., 2019). Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of lensing system in the universe (Li et al., 2019). Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution in this work.

### v4

@@ 2.4 Photometry
We obtained aperture photometry for SED fitting in each system by linearly summing the dereddened flux from pixels within the Einstein aperture in $g$-, $r$-, and $z$-band imaging data from DECaLS LS DR9. Since it is critical that we do not include light from the lensed source arcs as well as any interloping stars or galaxies that are not a member of the lensing cluster in the measured lens photometry, we created two masks to exclude non-cluster-member regions. The first mask removed flux from the lensed source arcs, and the second mask removed flux from interlopers as identified by a color and magnitude screening by eye. Accounting for masking, this linear integration was done for each of the 1000 bootstrapped Einstein apertures derived in Section 2.1. The resulting mean was adopted as the $g$-, $r$-, and $z$-band photometry for each system—with the standard deviation quantifying the geometric error. In quantifying the photometric error per-band, we first fit a sigma-clipped Gaussian to a histogram of the pixel values characterizing the background in each band for each system. The standard deviation of this Gaussian multiplied by $\sqrt{N}$, where $N$ is the number of pixels interior to the Einstein aperture, quantifies the statistical uncertainty in the photometric error for each band in each system. Next, the standard deviations for each Gaussian in a given photometric band across all systems were plotted as a histogram and fitted with another Gaussian. The standard deviation of this second Gaussian multiplied by $N$ characterizes the systematic uncertainty of the sky subtraction for each band across all systems. Taken together, the quadrature sum of the geometric Einstein aperture error, statistical photometric error, and systematic sky-subtraction error were adopted as the $g$-, $r$-, and $z$-band error for each system.
@@ 3.1 Total Mass
Following Remolina González et al. (2020), the enclosed total mass within the Einstein aperture was calculated using Equations (1) and (2)—where $D(z)$ refers to the angular diameter distance at redshift $z$—via a Monte Carlo approach. The enclosed total mass was computed 1000 times for each system, in which each calculation used a different Einstein radius, source redshift, and lens redshift randomly drawn from a Gaussian distribution with mean equal to the parameter value and standard deviation equal to the parameter error. Also described in Remolina González et al. (2020) is the need to apply an empirical correction described in Equation (3) based on the “completeness” of the lensed source arcs in each system—where $f(\theta_{E})$ is a cubic polynomial function specified in Table 1 of Remolina González et al. (2020). This is because the systematic bias and scatter of $M_{\Sigma}(<\theta_{E})$ has a dependence on the degree of symmetry for each system—in which larger-Einstein-radius systems have a stronger offset. The correction factor aims to remove this dependence as a function of Einstein radius in non-perfectly symmetric systems.
@@
For the purposes of applying this correction in each of the 177 strong lenses analyzed in this work, we measured the azimuthal coverage ($\phi$)—defined as the percentage of the Einstein aperture that was traced out by the lensed source arcs in a given system. For example, a lens with a tangential arc stretching from an azimuthal angle of $10^{\circ}$ to $100^{\circ}$ would have $\phi=0.25$. As was done by Remolina González et al. (2020), $\phi$ was used as an observable proxy for the degree of symmetry in each system to determine whether or not the correction factor was needed. Larger-Einstein-radius systems with a predominantly low $\phi$ tend to deviate from a symmetric dark matter halo and thus require the empirical correction. For spherically symmetric systems with a large $\phi$, the measured $M_{\Sigma}(<\theta_{E})$ is taken to be fairly unbiased, and thus the correction factor is not needed. We obtained $\phi$ by determining the fraction of the perimeter of the Einstein aperture that was subtended by the regions masking the lensed source arcs described in Section 2.4 in each system. Adopting the convention used by Remolina González et al. (2020) in their analysis, if $\phi<0.5$, Equation (3) was applied. If $\phi\geq 0.5$, it was not applied. Accounting for the empirical correction, the mean and standard deviation of the 1000 measurements of $M_{\Sigma}(<\theta_{E})$ were adopted as the enclosed total mass and its error, respectively, for each system.
@@ 3.2 Stellar Luminosity and Mass
We conducted parametric SED fitting using the photometry obtained in Section 2.4 with Prospector (Conroy et al., 2009; Conroy & Gunn, 2010; Johnson et al., 2021, 2023); five free parameters and three assumed parameters parameterized each SED. A summary of all eight parameters used in fitting can be found in Table 1. While the number of free parameters used in fitting is greater than the number of photometric band constraints for each system, parameters such as age, metallicity, and dust are correlated with one another (e.g., Worthey 1994; Thomas et al. 2017), and thus, SED fitting as done here remains sensible. Utilizing emcee (Foreman-Mackey et al., 2013) as implemented in Prospector with 84 walkers for a total of 6720 iterations, only the last 840 iterations for each of the 84 walkers (70,560 total parameter vectors) were taken as representing the posterior distribution for each of the five free parameters in each system in order to eliminate the significant burn-in sequence of the fitting process. After fitting, we adopted the 50th percentile of the posterior distribution as the value for each free parameter, and the greater difference between the 84th-50th percentile and 50th-16th percentile values as the error. Observed-frame SEDs for the systems shown in Figure 2 after SED fitting are shown in Figure 3.
Figure 4: A corner plot of the BCG redshift, enclosed total mass, and enclosed luminosity for the systems in this work plotted as black circles. Errors for each variable in each subplot are represented as red, blue, and green lines, respectively. A plane-of-best-fit is also shown in the top-right plot as well. This Figure is available as an animation. The animation depicts a rotating view of the plane and its associated constraints shown in the top right of the static Figure to emphasize its features in 3D. (An animation of this figure is available in the online article.)
To derive luminosity, we generated new SEDs with a random sample of 1024 free-parameter vectors from the posterior distributions of the best-fit SED in each system. We integrated over the rest-frame wavelength interval of 3000–7000Å, since this corresponds to the longest wavelength interval sampled by part of at least two bands of photometry across all BCG redshifts in this work (the lower bound of 3000Å gets redshifted out of the $\mathit{g}$ and into the $\mathit{r}$ band at high redshifts). This range also allows us to sample the rest-frame 4000Å break while recovering flux in longer wavelengths, which are significantly more luminous in cluster galaxies. After integrating the SEDs, we converted flux to luminosity in solar luminosities using the appropriate luminosity distance and adopted the mean and standard deviation from the 1024 random-parameter vectors as the luminosity and its error, respectively, for each system.
@@ 4 Results
As Figure 4 shows, the BCG redshift, enclosed total mass, and enclosed luminosity are correlated. By performing a linear regression scatter using emcee that accounts for errors in data on both axes (Hogg et al., 2010), we find the slope of the logarithmic correlation between the enclosed total mass and enclosed luminosity to be $0.528^{+0.022}_{-0.023}$ with an intrinsic scatter in luminosity of $0.115^{+0.008}_{-0.007}$ dex from the posterior distribution of the Markov Chain Monte Carlo (MCMC) fit. Indeed, we note that the most-massive clusters in this work also possess a large stellar mass and thus have a large luminosity. Also using emcee, we constrained a plane-of-best-fit taking into account the error in data on all three axes to describe this correlation—where X = $z_{{}_{BCG}}$, Y = $\mathrm{log}_{10}[M_{\Sigma}(<\theta_{E})/M_{\odot}]$, and Z = $\mathrm{log}_{10}[L(<\theta_{E})/L_{\odot}]$ with an intrinsic scatter on Z of $0.110^{+0.009}_{-0.008}$ dex—yielding:
$Z=0.148^{+0.051}_{-0.052}X+0.529^{+0.023}_{-0.022}Y+4.121^{+0.272}_{-0.298}$ (4)
The strong lensing candidates used to derive this relationship are, in our opinion, robust and unlikely to be non-lensing objects misclassified as bona fide lenses. Within the COOL-LAMPS candidate sample from which this subset derives, and indeed any other sample of strong lens candidates, there are a much larger number of less robust candidate systems. The identification of these systems as strong lenses may be more uncertain due to less obvious lensing geometry, and also—as is true in all flux-limited surveys—many of the notional lensed images will be only ambiguously detected. However, the simple and analytic relationship above, derived from a subsample of robust candidate strong lenses considering only quantities derived from the discovery survey observations, may have utility in differentiating bona fide strong lenses from non-lenses in the correspondingly larger samples of candidates where the evidence of lensing is not nearly as robust as the sample of candidates analyzed here. Moreover, this correlation offers a future opportunity to test simulations against observations—particularly those that aim to generate realistic galaxy populations in dense cluster environments.
@@
Figure 5: The total mass enclosed within the Einstein aperture for each system on the $x$-axis plotted against both the enclosed stellar mass and the enclosed stellar-to-total mass fraction on the $y$-axes. Points in Panels I and III are color coded according to the luminosity of the system in units of solar luminosities from stellar light enclosed within the Einstein aperture as measured within the rest-frame wavelength interval of 3000–7000Å. Points in Panels II and IV only represent systems in which multiple cluster members fall within the Einstein aperture. Linear regression lines in all four Panels are drawn from the posterior distributions of their respective MCMC.
@@ 5.1 Scaling Relations
In addition to the enclosed total mass and luminosity, we can also infer the observed enclosed stellar mass at the redshift of observation for each system. The stellar-mass posterior distribution from each SED constrained with Prospector describes the total amount of stellar mass formed over the lifetime of each system. However, the amount of stellar mass that we see at the redshift of observation is less than the total stellar mass formed over the lifetime of the system (e.g., such as supernovae or outflows; Li et al. 2017). To correct for this, Prospector also returns a surviving mass fraction at the conclusion of each fit that describes the percentage of the total stellar mass that remains in the system at the redshift of observation. By multiplying this percentage by the best-fit stellar mass formed, we obtain the stellar mass at the redshift of observation for each system. While we note that enclosed stellar mass is more complex to infer and subject to more significant systematics when compared to the more direct measures of enclosed total mass from strong-lensing geometry and luminosity, it nevertheless offers another point of comparison with those measurables.
@@ 5.1.1 Total Mass—Stellar Mass
In Panel I of Figure 5, we see a clearly positive correlation between the enclosed total mass and the enclosed stellar mass for all systems. By performing a linear regression with emcee that accounts for errors in data on both axes, we find that the slope of this correlation is $0.500^{+0.029}_{-0.031}$ with an intrinsic scatter of $0.104^{+0.005}_{-0.003}$ dex from the posterior distribution of the MCMC fit. We colored the points according to their luminosity to highlight how the intrinsic differences between individual systems contribute to the overall intrinsic scatter. In Table 2 of Kravtsov et al. (2018), they found $M_{500}$ (White, 2001) and the total stellar mass of the cluster within $R_{500}$ to be correlated with a logarithmic slope of $0.59\pm 0.08$—in agreement with the slope we derive.
Figure 6 shows a distribution of the physical scales within which mass and light was measured for this work and that of Kravtsov et al. (2018) and Gonzalez et al. (2013). It is striking that, despite differentiating between the small-scale apertures considered in this work as well as large-scale apertures with a physical distance scale up to 30 times larger, we measure an identical relationship across clusters in that the more-massive clusters are less efficient in converting baryons into stars as indicated by a sub-unity stellar-to-total mass slope. We caution against overinterpretation of this result given that, in both cases, the total and stellar masses are cylindrical masses projected through a 3D halo. Therefore, even a very small aperture samples some mass and galaxy light at large radii in projection. Nevertheless, this result suggests that cluster galaxies at a range of cluster-centric radii, at least those probed by the combination of our results and Kravtsov et al. (2018) and Gonzalez et al. (2013), follow similar efficiency dependencies as a function of total halo mass. In addition, since cluster galaxies at larger radii are located within larger apertures, yet retain the same slope between stellar and total mass, cluster galaxies in the core and the outskirts of the cluster (i.e. at a range of cluster-centric radii) must be similarly affected by the mass of the cluster at large in terms of the total baryon conversion efficiency. Panel II of Figure 5 further explores this regime, in which only those clusters where multiple cluster galaxies are contained within the Einstein aperture are plotted. We note that 103 of the clusters considered in this work met this criterion, and a linear regression for these systems gives a slope of $0.498^{+0.040}_{-0.039}$—identical to that of our entire sample.
Figure 6: A log-space histogram of the fitted Einstein radii for this work scaled to physical units (kpc) shown in blue. Also shown is a log-space histogram of $r_{500}$ for 21 galaxy clusters analyzed in Kravtsov et al. (2018) and Gonzalez et al. (2013).
Seeing as how the measured slopes agree across approximately two decades of cluster-centric radii suggests that whatever mechanism is responsible for suppressing baryon conversion efficiency is cluster wide and operating similarly regardless of the subpopulation of cluster member galaxies considered. What is unclear is whether this is predominantly a case of nature or nurture. Do the smaller halos that merge to form a massive cluster at later times have intrinsically lower baryon conversion efficiency, and that characteristic is inherited by the merged object (nature)? Or, is this suppression occurring at later times as a consequence of the halo growth through accretion and mergers (nurture)? That the stellar populations in cluster galaxies are much older than the assembled cluster-scale halo (e.g., Khullar et al. 2022) suggests the former.
5.1.2 Total Mass—Stellar-to-total-mass Fraction
@@
Andreon (2010) measured 52 clusters and found that the stellar mass fraction depended on $M_{200}$ with a slope of $-0.55\pm 0.08$. To compare directly with Andreon (2010), we plot the enclosed stellar-to-total mass fraction against the enclosed total mass in Panel III of Figure 5 and see a clearly negative correlation (again with the same color-mapping as Panel I to emphasize intrinsic scatter). Also using emcee, we find that this correlation has a slope of $-0.495^{+0.032}_{-0.033}$ and an intrinsic scatter of $0.103^{+0.005}_{-0.002}$ dex. Categorizing in Panel IV identically to Panel II, we find that the multiple-galaxy systems are correlated with a slope of $-0.497^{+0.043}_{-0.047}$ and intrinsic scatter of $0.105^{+0.007}_{-0.004}$ dex. We again find that the slope for the multiple-galaxy systems is statistically identical to the slope for the entire sample shown in Panel III. The slopes we derive in Panels III and IV are statistically indistinguishable from those of Andreon (2010), who measured out to the physical distance scale of $R_{200}$—systematically larger than $R_{500}$ used in Kravtsov et al. (2018) and Gonzalez et al. (2013) and much larger than this work.
While a later paper by Andreon (2012) found a shallower slope of $-0.38\pm 0.07$ with a much smaller sample of only 12 galaxy clusters, the slope we derive with 177 clusters and that of Andreon (2010) supports the prescription of a slope of $\sim-0.49$ from Bode et al. (2009) to reproduce observed X-ray scaling relations without any supplemental feedback from nonstellar sources. Kravtsov et al. (2018) also noted that the steepness of the slope they derived, in agreement with our own, indicates weaker feedback than cosmological simulations (e.g., Kravtsov et al. 2005; Martizzi et al. 2014) suggest. While this conclusion was not explicitly probed in this work, the large difference between the aperture sizes we use and those of Andreon (2010) and Kravtsov et al. (2018) and the similarity of the derived slopes is suggestive of scale invariance with respect to feedback strength in addition to baryon conversion efficiency across a wide range of cluster-centric radii.
@@ 5.2 Photometric Bias and Error
Deriving light in Section 2.4 via directly integrating over the Einstein aperture presents some challenges. Since we identified red-sequence cluster members by using observed color as a proxy for cluster membership, it is possible that we may have included intrinsically bluer galaxies at redshift $z>z_{{}_{BCG}}$ that appear to be the same color as the cluster members. Conversely, we may have excluded intrinsically bluer bona fide cluster members. However, star-forming cluster members have less stellar mass and are intrinsically rarer in cluster cores (e.g., Dressler et al. 1997, 2004), and we do not expect that these missed cluster galaxies introduce a significant loss to the measured cluster light. In addition, bluer galaxies at redshift $z>z_{{}_{BCG}}$ appear relatively fainter, which are also unlikely to bias our methodology of measuring cluster light. Interlopers such as foreground stars that appear in front of cluster galaxies are forced to be masked, but any underlying cluster galaxy flux is typically minimal due to the limited angular size of such stars.
@@
As shown in Figure 1 and quantified in Section 2.1, our sample becomes incomplete at a measured Einstein radius of $3.09^{\prime\prime}$—falling to 50% incomplete by $2.11^{\prime\prime}$. To some extent, this is a consequence of our chosen sample of cluster-scale lenses, which we would generally expect to have larger Einstein radii. However, we caution that there exists a large scatter between the total halo mass and Einstein radius (Fox et al., 2022). Systems whose arcs are separated from the central galaxy at small radii may also be located directly within the discernible BCG light. This would not only make them hard to visually notice, but it would also make them unsuitable for this analysis because the source light would be significantly blended together with the lensing cluster light. While we thus expect and demonstrate that this work undersamples the small-Einstein-radius regime, it is not apparent that this significantly alters our findings.
In addition, our use of relatively shallow ground-based imaging limits the $5\sigma$ point-source magnitude depth of objects in the $g$-, $r$-, and $z$-band filters to $\approx$ 24.7, 23.9, and 23.0, respectively in the average case of two exposures per filter (Dey et al., 2019). Systems with a lens redshift $z\gtrapprox 1$ are barely visible even in $\mathit{z}$-band LS DR9 imaging data, and they are also simply a rarer type of strong lens in the Universe (Li et al., 2019). Both the data depth and intrinsic redshift distribution of lenses shape the redshift distribution of the clusters in this work.